Reemu read (1/5)th pages of a book. If she reads further 40 pages, she would have read (7/10)th pages of the book. How many pages are left to be read?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question it is given that,

Reemu read (1/5)thpages of a book.
Let us assume the total number of pages in the book be ‘P’.Then, number of pages read by Reemu = (1/5) of P

= (1/5) × P
And also it is given in the question, If she reads further 40 pages, she would have read

(7/10)thpages of the book. = (7/10) × P So,

((1/5) × P) + 40 = (7/10) × P
(P + 200)/5 = (7P/10)
By cross mutliplication we get,

2P + 400 = 7P 5P = 400
P = 400/5 ... [divide both numerator and denominator by 5]
P = 80
Then, pages read by Reemu = Total pages–Pages read
P–(7P/10) = (3P/10)

(3/10) × 80

24 pages
Therefore, 24 pages are left to be read

Thanks